Tess is 4 years older than her sister, Ana. In 4 years,she will be 37 years more than 1/4 as old as Ana will then be. How old are the girls now?

Answer by mananth(16946)

Tess is 4 years older than her sister, Ana. In 4 years,she will be 37 years more than 1/4 as old as Ana will then be. How old are the girls now?
Ana = x years
Tess = x+4
...
after 4 years
Ana = x+4
Tess = x+4+4 = x+8 years
...
x+8 = 37 + 1/4 * (x+4)
x-29=1/4(x+4)
multiply by 4
4(x-29)=x+4
4x-116=x+4
4x-x=116+4
3x=120
/3
x= 40 years Ana's age
Tess = x+4 = 44 years
..
m.ananth@hotmail.ca